OpenStudy (king):

A stone is thrown vertically upwards with initial velocity u.Find the distance travelled by it in time 1.5u/g

OpenStudy (king):

i got 3u^2/8g but answer given is 5u^2/8g

OpenStudy (ash2326):

s=ut+1/2at^2 u=u a=-g t=1.5u/g s=u*1.5u/g-1/2*g*2.25*u^2/g^2 s=1.5u^2/g-2.25u^2/2g s=0.75u^2/2g s=0.375u^2.g s=3u^2/8g

OpenStudy (anonymous):

When s=u*1.5u/g-1/2*g*2.25*u^2/g^2 what happened to the -1/2*g*2.25*u^2/g^2 to get -2.25u^2/2g ?